logbase3(logbaseX(logbase416))=-1
I thought perhaps I must simply just add (which multiplies) so i took it out of the brackets.
I was getting
logbase3 +logbaseX +logbase4 4^2 <-fours cancel out right?
Then I got logbase3 +logbase X +2=-1
then log base3 +log base X=-3
Then I am lost.
Please help?

Answer by stanbon(75887)

logbase3(logbaseX(logbase4(16))=-1
-------
log3(logx(log4(16)) = -1
----
Work from right to left:
---
= log3(logx(2)) = -1
---
logx(2) = 3^-1
----
logx(2) = 1/3
---
2 = x^(1/3)
x = 2^3 = 8
==================
Cheers,
Stan H.